OpenStudy (cruze):

Check my answers please :) A fighter jet landing on an aircraft carrier's flight deck that has a length of 300.0 m must reduce it's speed from about 42.5 m/s to 0 m/s in 2.0 s. What is the jet's acceleration? I got 44.5 m/s2 (45 m/s2 when using sig figs). Is my answer true?

OpenStudy (cruze):

I actually did my work wrong and got an answer of -21.25 m/s2 (-21 m/s2 when using sig figs).

OpenStudy (cruze):

Is it still correct?

OpenStudy (jiteshmeghwal9):

Initial speed(u)=42.5 m/s Final speed(v)=0m/s Time taken (t)=2 s Acceleration=(v-u)/t Acceleration=(0-42.5)/2

OpenStudy (jiteshmeghwal9):

-21m/s^2 should be correct